Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using anReq 1A Reg 1B Reg 2A Reg 2B Req3 For direct materials, the materials were purchased from a new supplier who is anxious to entReq 1A Req 1B Reg 2A Reg 2B Reg 3 For direct labor, compute the rate and efficiency variances. (Indicate the effect of each vReq 1A Reg 1B Req 2A Reg 2B Reg 3 In the past, the 26 technicians employed in the production of Fludex consisted of 4 seniorReq 1A Reg 1B Req 2A Req 2B Req3 Compute the variable overhead rate and efficiency variances. (Indicate the effect of each va
Answer #1

Part 1 a

Standard price

-

Actual price

X

Actual quantity

=

Variance

Materials price variance

24

22.60

12500

17500

F

Standard quantity

-

Actual quantity

X

Standard price

=

Variance

Material quantity variance

9430

9500

24

1680

U

Actual price = 282500/12500

Actual quantity used = 12500-3000 = 9500

Standard quantity = 4100*2.30 = 9430

Part 1 b

Yes, the contract probably should be signed. The new price of $22.60 per ounce is substantially lower than the old price of $24 per ounce, resulting in a favorable price variance of $17500 for the month. Moreover, there is a little or no problem in production which is depicted from the small materials quantity variance for the month.

Part 2 a

Standard rate

-

Actual rate

X

Actual hours

=

Variance

Labor rate variance

15.0

14.0

3900

3900

F

Standard hours

-

Actual hours

X

Standard rate

=

Variance

Labor efficiency variance

3280

3900

15

9300

U

Actual hours = 26*150 = 3900

Standard hours =4100*0.8 =3280

Part 2 b

No, it is not advisable to continue new labor mix probably. In spite of decrease in the average hourly labor cost from $15 to $14, which results into the $3900 favorable labor rate variance. The overall labor efficiency variance is greater than the labor rate variance. This shows that the new labor mix increases overall labor costs.

Part 3 a

Standard rate

-

Actual rate

X

Actual hours

=

Variance

Variable overhead rate variance

3.50

2.05

3900

5655

F

Standard hours

-

Actual hours

X

Standard rate

=

Variance

Variable overhead efficiency variance

3280

3900

3.50

2170

U

Actual rate per unit = 8000/3900 = 2.05